OpenStudy (anonymous):

A 20% acid solution is mixed with a 70% acid solution to get 50 liters of a 40% solution. How much of the 70% solution is used? 20 liters 25 liters 30 liters

OpenStudy (anonymous):

\[ x+y=50\\ 0.20x+0.70y = 0.40(x+y) \]

OpenStudy (anonymous):

If you solve for \(y\), you will get the answer.

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

Does the equation make sense?

OpenStudy (anonymous):

i tried that but for the second equation i did .2x+.7y=.4 and thats all so thats what i did wrong and yeah it makes sense thanks so much

OpenStudy (whpalmer4):

yep, need to think of the second equation in terms of the amount of stuff...0.2 liters of acid in each of x liters of 20% solution + 0.7 liters of acid in each of y liters of 70% solution = 0.4 liters of acid in each of 50 liters of 40% solution

OpenStudy (anonymous):

30 litres

OpenStudy (whpalmer4):

check your answer: adding 30 liters of 70% acid to 20 liters of 20% acid: 30*0.7 + 20*0.2 = 21+4=25 liters of acid in 50 liters of solution or a 50% concentration. You want a 40% concentration.

OpenStudy (anonymous):

sorry 20 liters

OpenStudy (anonymous):

other one is 30..

OpenStudy (whpalmer4):

30*.2 + 20*.7 = 6 + 14 = 20 liters of acid in 50 liters of solution = 40%, that's correct.

OpenStudy (whpalmer4):

heartbreaking to do the problem right and get it wrong because you mislabeled or misinterpreted the results!

OpenStudy (anonymous):

lol.. happens with me all the times

OpenStudy (anonymous):

thanks